OpenStudy (anonymous):

Lila earns $8.75/h at her job. How many hours must Lila work to earn $105? Complete the equation to model the problem. Use h to represent the unknown value.

OpenStudy (anonymous):

@Kit_Kat_Nat_<3 @KissMyAxe @marylou004

OpenStudy (anonymous):

8.75h=105

OpenStudy (anonymous):

it wont let me type in numbers

OpenStudy (anonymous):

three more plz

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Which equations model this problem? A recipe that makes 1 serving calls for 0.25 L milk. How many servings can Rohit make if he uses 2 L of milk? Use s for the number of servings. Choose all answers that are correct. A. B. C. D.

OpenStudy (anonymous):

what are the choices they're blank

OpenStudy (anonymous):

Please put this under math

OpenStudy (anonymous):

A. 0.25/2=s

OpenStudy (anonymous):

B.0.25*s=2

OpenStudy (anonymous):

C.2/0.25=s

OpenStudy (anonymous):

D. s/0.25=2

OpenStudy (anonymous):

@Asuna-Yuuki

OpenStudy (anonymous):

i think its B and C

OpenStudy (anonymous):

The temperature of a gas decreases 0.045°C every minute for 8 min. What is the change in the temperature of the gas? Use a positive or negative number to express the change in the temperature of the gas.

OpenStudy (anonymous):

do you have to write an equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-0.045

OpenStudy (anonymous):

A diver descends below the surface of the water at a rate of 9.3 m/min. What is the diver’s altitude relative to sea level after 3 min? m

OpenStudy (anonymous):

@Asuna-Yuuki

OpenStudy (anonymous):

-27.9

OpenStudy (anonymous):

(1 pt) Would each situation be represented by a positive or a negative number? Column A Column B 1. Three withdrawals of n dollars are made from a bank account that started with a $0 balance. Express the balance of the bank account. 2. Every hour for h hours the temperature goes down 0.2 degrees. Express the change in temperature. 3. A tank of gasoline is 0.75 full. The tank holds t gallons. Express how much gas is in the tank. 4. A hot air balloon starts at ground level and ascends at a rate of 10 ft per s. Express the change in altitude. A. positive B. negative

OpenStudy (anonymous):

Number 1 you would get a negative number because withdraw is take out and he has $0 so all you're doing is taking away from 0 which would give you a negative number

OpenStudy (arabpride):

\(\Large\color{purple}{Hey!}\) \(\normalsize\color{purple}{Glad~you~got~help~but~please~ask~in~the~correct~section~next~time:}\) \(\Large\color{green}{Mathematics~ :D}\)